# Planar percolation and the loop $\mathrm{O}(n)$ model

### v1

@@ 1.1. Site percolation models on planar graphs
In the case where $G$ can be embedded in the plane with no accumulation points, this conjecture has been confirmed by Haslegrave and Panagiotis [HP21] ($p_{c}<1/2$) and Li [Li23] ($p_{u}\geq 1-p_{c}$). We expect that the minimal degree $7$ bound is not necessary for the bound $p_{u}\geq 1-p_{c}$. Indeed, our result implies that
@@ 1.2. Corollaries for percolation on unimodular invariantly amenable graphs
A bond percolation on a graph $G=(V,E)$ is a random variable $\omega$ that takes values in $\{0,1\}^{E}$. We identify $\omega$ with the spanning subgraph of $G$ given by $\{\omega=1\}$; we say that it is finitary if all connected components are finite, almost surely, and invariant if $(G,\rho,\omega)$ is unimodular in a sense that (MTP) holds also when $f$ depends on both $(G,\rho)$ and $\omega$ (see [AHNR18, Section 2] for more details). For $v\in V$, define $K_{\omega}(v)$ as the connected component of $v$ in $\omega$; for $U\subseteq V$, define $\partial U$ as the set of vertices in $U$ that are adjacent to $V\smallsetminus U$. We say that $(G,\rho)$ is invariantly amenable if
@@
As was shown in [AHNR18], the assumptions of the corollary above are known to be equivalent to many other properties of ‘parabolic’ planar unimodular random maps. One such condition is that $(G,\rho)$ can be expressed as the Benjamini–Schramm (or local) limit of finite planar graphs. In this setting, Peled [Pel20] proved the existence of a uniform lower bound on $p_{c}(G)$, and showed the circle packing of such graphs exhibits exponential decay below that threshold. Our bound is optimal since the site percolation on the triangular lattice has $p_{c}=1/2$ [Kes80].
The corollary follows from Theorem 1 once one rules out infinitely many infinite components. We do this by extending the approach of Burton and Keane to invariantly amenable unimodular graphs. In fact, our approach extends beyond Bernoulli percolation to the framework of divide and color models (Corollary 1.2).
For a graph $G=(V,E)$, we define $\mathrm{ER}(V)$ as the set of all equivalence relations on $V$. Given $\mu$ a probability measure on $\mathrm{ER}(V)$, a divide and color model of parameter $p\in[0,1]$ is a site percolation $\sigma$ on $G$ defined as follows. First, one samples a partition $P=\{P_{i}\}_{i=1}^{\infty}$ from $\mu$. Then, for each $i\in\mathbb{N}$, we assign all vertices of $P_{i}$ the value $\sigma=1$ with probability $p$ and $\sigma=0$ otherwise, independently of all other classes. The model was first introduced by Häggström [Häg01] in the case when $\mu$ is Bernoulli bond percolation; the critical value function in this case was studied in [BBT13b, BBT13a]. The Gibbs properties of the measure when $\mu$ is the random-cluster model was considered in [Bál10].
@@ 1.4. Sketch of the proof of Theorem 1
The main step of the proof is splitting $\partial\Omega_{n}$ into two arcs, each connected to infinity in $\Omega_{n}^{c}$ with probability almost one. Denote the vertices on $\partial\Omega_{n}$ by $v_{1},\dots,v_{L}$, taken in clockwise order. For each $i,j\in\{1,\dots,L\}$, let $p_{i,j}$ denote the probability that $\{v_{i},v_{i+1},\dots,v_{j}\}$ does not connect to infinity in $\Omega_{n}^{c}$. Clearly, $p_{1,i}$ is decreasing and $p_{1,L}<\varepsilon$. Define
@@
By the minimality of $i^{*}$, we have that $p_{1,i^{*}-1}\geq 2\sqrt{\varepsilon}$, whence $p_{i^{*},L}<\sqrt{\varepsilon/2}$. Since $\sigma$ is dominated by $1-\sigma$, we have that
@@ Acknowledgements
We thank Diederik van Engelenburg for insightful discussions on unimodular graphs. The project started from an open problem session at a conference at the ETH Zurich in 2024, and we thank Hugo Duminil-Copin and Vincent Tassion for organizing this wonderful event. Parts of the project were completed during the visits of AG and MH to Northeastern University and the University of Innsbruck, and we would like to thank these institutions for their hospitality.
@@ 2. Planar percolation models: proof of Theorem 1
Both Corollaries 1.1 and 1.2 follow from Theorem 1 if one can rule out the existence of infinitely many infinite connected components. A beautiful argument of Burton and Keane [BK89] treats a large class of bond percolations on transitive amenable graphs. Here we adapt it to the setting of site percolations on unimodular invariantly amenable graphs. Note that a site percolation might not have any trifurcation point due to combinatorial constraints (e.g. the Kagome lattice). In order to resolve this issue, we sample an auxiliary uniform spanning forest $\tau$ on top of our site percolation $\sigma$ and show that $\tau$ does have trifurcation points. We then extend the argument to unimodular graphs using (MTP) in a fairly standard way. We provide all details of the proof of Corollary 1.1 and then explain how to modify it to the more general setting of Corollary 1.2.
@@
For $n\in\mathbb{N}$, consider the subgraph of $G$ induced by $\{\sigma=1\}\cap\Lambda_{n}(\rho)$ and denote its connected component containing $\rho$ by $\mathcal{C}_{n,\rho}$. Define $\mathsf{UST}_{n,G,\sigma}$ as the uniform spanning tree measure on $\mathcal{C}_{n,\rho}$. It is standard that the UST satisfies a negative correlation inequality, whence the limit of $\mathsf{UST}_{n,G,\sigma}$ as $n$ tends to infinity exists; see [BLPS01]. The limit is called free uniform spanning forest, we denote it by $\mathsf{FUSF}_{G,\sigma}$. A more recent work [AHNR18] shows that, on unimodular invariantly amenable graphs, the FUSF is connected, almost surely. To keep our presentation self-contained, we will not rely on this result.

### v2

@@ 1.1. Site percolation models on planar graphs
In the case where $G$ can be embedded in the plane with no accumulation points, this conjecture has been confirmed by Haslegrave and Panagiotis [HP21] ($p_{c}<1/2$) and Li [Li23] ($p_{u}\geq 1-p_{c}$). We expect that $p_{u}\geq 1-p_{c}$ holds also without any assumptions on the degrees of the graph. Indeed, our result implies that
@@ 1.2. Corollaries for percolation on unimodular invariantly amenable graphs
A bond percolation on a graph $G=(V,E)$ is a random variable $\omega$ that takes values in $\{0,1\}^{E}$. We identify $\omega$ with the spanning subgraph of $G$ given by $\{\omega=1\}$; we say that it is finitary if all connected components are finite, almost surely, and invariant if $(G,\rho,\omega)$ is unimodular in a sense that (MTP) holds also when $f$ depends on both $(G,\rho)$ and $\omega$ (see [AHNR18, Section 2] for more details). For $v\in V$, define $K_{\omega}(v)$ as the connected component of $v$ in $\omega$; for $U\subseteq V$, define $\partial U$ as the set of vertices in $U$ that are adjacent to $V\smallsetminus U$. We say that $(G,\rho)$ is (vertex) invariantly amenable if
@@
Our definition via inner vertex boundaries agrees with [AL07] and is different from [AHNR18], which uses edge boundaries. Note that if a graph is edge invariantly amenable in the sense of [AHNR18], then it is also vertex invariantly amenable. Moreover, when the root has a finite expected degree, being unimodular and (edge) invariantly amenable is equivalent to being the Benjamini–Schramm (or local) limit of finite planar maps [LT80, AHNR18].
@@
Previously, Peled [Pel20] proved the existence of some $c>0$ such that $p_{c}>c$ for the Benjamini–Schramm limit of arbitrary finite planar maps; that work also shows that the circle packing of such graphs exhibits exponential decay at all $p<c$. We do not know whether all such graph are (vertex) invariantly amenable, but, as mentioned above, the implication is true when the root has a finite expected degree. Corollary 1.1 thus gives that $p_{c}(G)\geq 1/2$ if $(G,\rho)$ is the Benjamini–Schramm limit of finite planar maps and $\rho$ has a finite expected degree. This bound is optimal since the site percolation on the triangular lattice has $p_{c}=1/2$ [Kes80].
We further extend the statement to the framework of divide and color models (Corollary 1.2). For a graph $G=(V,E)$, we define $\mathrm{ER}(V)$ as the set of all equivalence relations on $V$. Given $\mu$ a probability measure on $\mathrm{ER}(V)$, a divide and color model of parameter $p\in[0,1]$ is a site percolation $\sigma$ on $G$ defined as follows. First, one samples a partition $P=\{P_{i}\}_{i=1}^{\infty}$ from $\mu$. Then, for each $i\in\mathbb{N}$, we assign all vertices of $P_{i}$ the value $\sigma=1$ with probability $p$ and $\sigma=0$ otherwise, independently of all other classes. The model was first introduced by Häggström [Häg01] in the case when $\mu$ is Bernoulli bond percolation; the critical value function in this case was studied in [BBT13b, BBT13a]. The Gibbs properties of the measure when $\mu$ is the random-cluster model was considered in [Bál10].
@@
Corollaries 1.1 and 1.2 follow from Theorem 1 once one rules out infinitely many infinite components. The latter statement was proven in a number of cases: the classical argument of Burton and Keane [BK89] treats shift invariant site percolation processes on $\mathbb{Z}^{d}$ with the finite energy property. This argument was then extended in [AL07, Sections $6,8$] based on [LS99, BLPS99, BLPS01] (see also [AHNR18, Theorem 5.9]) to insertion-tolerant invariant bond percolation processes on unimodular graphs. In [LP17, Theorem 7.9], insertion-tolerant invariant site and bond percolation processes on transitive graphs were treated. The arguments in these works use insertion-tolerance only to construct an infinite component with at least three ends, so the assumption can be weakened. The combination of these arguments extends the statement to the setting of Corollaries 1.1 and 1.2; we provide short proofs for completeness.
@@ 1.4. Sketch of the proof of Theorem 1
The main step of the proof is splitting $\partial\Omega_{n}$ into two arcs, each connected to infinity in $\Omega_{n}^{c}$ with probability almost one (a similar idea was also used in [Tas16]). Denote the vertices on $\partial\Omega_{n}$ by $v_{1},\dots,v_{L}$, taken in clockwise order. For each $i,j\in\{1,\dots,L\}$, let $p_{i,j}$ denote the probability that $\{v_{i},v_{i+1},\dots,v_{j}\}$ does not connect to infinity in $\Omega_{n}^{c}$. Clearly, $p_{1,i}$ is decreasing and $p_{1,L}<\varepsilon$. Define
@@
By the minimality of $i^{*}$, we have that $p_{1,i^{*}-1}\geq\sqrt{2\varepsilon}$, whence $p_{i^{*},L}<\sqrt{\varepsilon/2}$. Since $\sigma$ is dominated by $1-\sigma$, we have that
@@ Acknowledgements
We thank Diederik van Engelenburg, Russell Lyons, and Gabor Pete for insightful discussions on unimodular graphs. The project started from an open problem session at a conference at the ETH Zurich in 2024, and we thank Hugo Duminil-Copin and Vincent Tassion for organizing this wonderful event. Parts of the project were completed during the visits of AG and MH to Northeastern University and the University of Innsbruck, and we would like to thank these institutions for their hospitality.
@@ 2. Planar percolation models: proof of Theorem 1
Both Corollaries 1.1 and 1.2 follow from Theorem 1 if one can rule out the existence of infinitely many infinite connected components. A beautiful argument of Burton and Keane [BK89] treats a large class of bond percolations on transitive amenable graphs. Here we adapt it to the setting of site percolations on unimodular invariantly amenable graphs. Note that a site percolation might not have any trifurcation point due to combinatorial constraints (e.g. the Kagome lattice). In order to resolve this issue, we sample an auxiliary uniform spanning forest $\tau$ on top of our site percolation $\sigma$ and show that $\tau$ does have trifurcation points (this is reminiscent to [LP17, Lemma $7.7$]). We then extend the argument to unimodular graphs using (MTP) in a fairly standard way. We provide all details of the proof of Corollary 1.1 and then explain how to modify it to the more general setting of Corollary 1.2.
@@
For $n\in\mathbb{N}$, consider the subgraph of $G$ induced by $\{\sigma=1\}\cap\Lambda_{n}(\rho)$ and denote its connected component containing $\rho$ by $\mathcal{C}_{n,\rho}$. Define $\mathsf{UST}_{n,G,\sigma}$ as the uniform spanning tree measure on $\mathcal{C}_{n,\rho}$. It is standard that the UST satisfies a negative correlation inequality, whence the limit of $\mathsf{UST}_{n,G,\sigma}$ as $n$ tends to infinity exists; see [BLPS01]. The limit is called free uniform spanning forest, we denote it by $\mathsf{FUSF}_{G,\sigma}$. Below we do not rely on [AHNR18, Theorem $5.13$] establishing that the FUSF is connected a.s. since that work assumes that $\rho$ has a finite expected degree.